OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? A) y = 2x + 1 B) y = 2x + 3 C) y = 2x - 8 D) y = 2x - 3

OpenStudy (anonymous):

y - 5 = 2(x - 2) y - 5 = 2x - 4 y = 2x + 1

jimthompson5910 (jim_thompson5910):

y = mx+b 5 = (2)(2)+b 5 = 4+b 5-4 = b 1 = b b = 1 So the equation of the line is y = 2x+1

OpenStudy (anonymous):

For this type of question, plug in x( value of 2 in this case ) into the equation and see if the y comes out as 5. Then you look at the slope by seeing the coefficient of x, and if it's 2. A satisfies both conditions, and so it's the answer.

OpenStudy (anonymous):

Thank y'all<3